OpenStudy (anonymous):

derive cos^2(x)+sin(x). Then find the critical numbers.

OpenStudy (anonymous):

Derive or Derivate ?

OpenStudy (anonymous):

Sorry, derivative. I had a late night and just stumbled upon this website.

OpenStudy (anonymous):

\[f'(x)=2*\cos(x)*-\sin(x)+\cos(x)\]

OpenStudy (anonymous):

Wait, how? Thanks though...

OpenStudy (anonymous):

do you know the chain rule?

OpenStudy (anonymous):

Yes, I got {(-sinx*cosx)+(cosx*-sinx)} + cosx.

OpenStudy (anonymous):

what you calculated is the same as what I wrote. :)

OpenStudy (phi):

yes

OpenStudy (phi):

btw, did you use the product rule on cos^2(x) i.e. write it as cos(x) * cos(x) ? you could also use the "power rule" d u^2 = 2 u du or in this case d cos^2 x = 2 cos(x) d cos(x) = 2 cos(x) (-sin x) or -2 cos(x) sin(x)

OpenStudy (anonymous):

But I don't know how to solve once set to zero..

OpenStudy (anonymous):

hint: \[Cos(x)*(-2*Sin(x)+1)=0\]

OpenStudy (phi):

-2sinx cosx+ cosx= 0 if we used a= sin x and b= cos x (just renaming ) this is the same problem as -2 ab + b =0 factor out b: b(-2a+1) = 0 either b is 0 or -2a+1 is 0 or, using the originals cos x = 0 or -2 sin x + 1 = 0

OpenStudy (phi):

cos x is zero at pi/2 , pi/2+ pi, pi/2 + 2pi , etc in general \[ \frac{\pi}{2} + n \pi \]

OpenStudy (anonymous):

so the pi's would cancel, and 2 would be my answer?

OpenStudy (phi):

no. the cosine "goes forever" and it crosses the x-axis many times http://www.mathsisfun.com/algebra/trig-sin-cos-tan-graphs.html the "zeros" occur at pi/2 (90 degrees) and repeat every pi radians (every 180 deg) in both directions.